# Perceptual Decoupling for Scalable Multi-modal Reasoning via Reward-Optimized Captioning

### v1

@@ Results.
Sol. Cap. LLM CRO Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
31.5 36.0 10.0 39.2 11.4 3.8 22.8 22.1
@@
64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
60.1 36.3 30.2 47.6 24.4 14.6 44.5 36.8
59.5 36.2 12.7 47.8 25.0 13.4 35.1 32.8
63.8 34.5 30.3 51.7 27.4 17.4 43.9 38.4
64.5 36.3 39.3 51.7 31.5 21.4 46.5 41.6
@@
70.3 25.8 41.0 57.3 34.4 19.4 46.1 42.0
68.3 38.0 41.0 53.9 31.2 22.4 47.4 43.2
54.4 35.8 27.9 50.6 23.7 7.6 34.9 33.6
70.9 37.6 42.3 56.4 37.3 24.6 48.3 45.3
71.9 38.9 45.3 58.7 37.1 27.0 49.7 46.9
4.2 Ablation Study
In this section, we first investigate the effectiveness of the proposed components, i.e., Reasoning-Aligned Perceptual Decoupling and Caption Reward Optimization. Then, we ablate the choices on reward computation during training. We use the same training configurations as in Sec. 4.1. Qwen2.5-VL-3B/7B are adopted for ablations due to resource constraints. Unless otherwise specified, we use R1-Distilled-7B as the default LLM reasoner throughout the training and inference stages.
Reasoning-Aligned Perceptual Decoupling & Caption Reward Optimization.
For simplicity, we use the following shorthand markers to indicate the components:
  • LLM indicates the usage of the text-only LLM reasoner $M_{R}$.
  • Cap. denotes the use of the query-conditioned captions generated by the MLLM $M_{V}$.
  • Sol. denotes the (tentative) solution provided by the MLLM $M_{V}$. Note that if the reasoner $M_{R}$ (i.e., LLM above) is not used, it is considered as the final output solution.
  • CRO denotes the use of Caption Reward Optimization that optimizes the quality of the generated captions.
We assess the RACRO framework by progressively enabling its components, as in Table 2. Starting from the baseline MLLM, we first introduce the tentative solution as the input to the LLM reasoner. This alone leads to notable performance improvements—e.g., from 34.6% to 36.8% for Qwen2.5-VL-3B and from 42.0% to 43.2% for the 7B model—demonstrating the reasoner’s ability to refine intermediate outputs from the MLLM. However, on tasks such as MathVista, MMMU, and WeMath, performance declines after introducing the LLM with tentative solutions only. To compensate for the lack of visual details, we incorporate query-conditioned captions alongside the tentative solution, which effectively restores the performance. Surprisingly, when captions are adopted alone without tentative solutions, benchmark performance drops significantly, suggesting that while captions can provide valuable visual context, they are often noisy—suffering from hallucinations and missing critical details. In contrast, the tentative solution acts as a reliable fallback for the reasoner when the visual signal is weak or ambiguous. To improve the quality of the query-conditioned captions, we apply the proposed Caption Reward Optimization strategy. As a result, we observe consistent performance gains across all tasks, confirming the quality issue of the captions and the effectiveness of the Caption Reward Optimization strategy.
Table 3: Ablation studies on (a) reasoner types and (b) inputs for reward computation.
Configuration Q-VL-3B Q-VL-7B
w/o CRO (baseline) 38.4 45.3
(a) Reasoner Types
Qwen-2.5-7B-Instruct 40.4 45.1
R1-Distilled-1.5B 40.9 45.2
R1-Distilled-7B 41.6 46.9
(b) Reasoner Inputs
Cap. & Sol. 40.8 45.8
Cap. 41.6 46.9
Choices on Reward Computation.
Conducting CRO requires computing the reward of a caption rollout, which involves running the reasoner LLM. Here we discuss (i) what type of reasoner and (ii) what kind of inputs to the reasoner lead to more accurate reward computation. Specifically, for reasoners, we consider R1-Distilled-1.5B/7B and Qwen-2.5-7B-Instruct. For the reasoner input, we consider “Cap.”, which only includes the generated caption and “Cap. & Sol.” that further includes the MLLM’s tentative solution (pre-generated before training). Note that the later is exactly the configuration adopted in the proposed Perceptual Decouple framework.
Table 3 shows the average performance with Qwen2.5-VL series as MLLMs (i.e., Q-VL-3B/7B) under different reasoner types and inputs. We observe that using smaller reasoners (i.e., R1-Distilled-1.5B) or models less specialized in reasoning (i.e., Qwen-2.5-Instruct) leads to worse overall performance. Intuitively, models performing poorly in reasoning are prune to generate incorrect solutions even though captions are accurate and comprehensive, providing noisy reward signals for CRO. However, larger reasoners significantly slow down the reward computing. Thus, we adopt R1-Distilled-7B as the reasoner as it strikes a balance between computational efficiency and reward accuracy by default.
Regarding the input to the reasoner, we found passing only captions yields the best performance. Although “Cap. & Sol.” perfectly aligns with the Perceptual Decoupling framework during inference time, including a tentative response, allows the reasoner to take a shortcut that ignore the captions during reasoning, thereby producing a noisy reward signal for the captions.
4.3 Analysis
Scalability of CRO.
In a practical scenario, after conducting CRO on an MLLM with a certain LLM (e.g., R1-Distilled-7B), it is desirable to switch to new LLMs (e.g., QwQ-32B) without additional training. We refer to this as the scalability of CRO and reveal that CRO indeed possesses this property.
Figure 6 shows the overall task performance under the Perceptual Decoupling framework using different reasoners across varying model sizes and generations. The MLLMs adopted are from the Qwen2.5-VL series, both with and without CRO. As shown, although all MLLMs were optimized with CRO using R1-Distilled-7B, switching to different LLMs—whether in size (R1-Distilled-1.5B/32B) or generation (QwQ-32B)—still results in a notable performance boost over their non-CRO counterparts. This confirms the scalability of CRO. Furthermore, we observe that this performance gain becomes more pronounced with larger or stronger LLMs. This is because, without CRO, these models are significantly bottlenecked by inaccurate or low-quality visual captions, preventing them from leveraging their full reasoning potential. In contrast, smaller or weaker LLMs are limited not only by caption quality but also by their own reasoning capacity, making the relative improvement from CRO less substantial. This highlights the effectiveness of CRO in improving the quality of image captions. Further, with stronger LLMs, the performance gap between 3B and 7B gradually diminishes. This reveals that reasoning is better decoupled from perception.
Figure 5: Ablation on the choice of reasoner.
Figure 6: Ablation on CRO and GRPO.
CRO and GRPO are Complementary.
The proposed CRO objective effectively enhances the quality of query-conditioned captions, thereby improving overall performance. Another direction for improving the reasoning capability of the LLM is to produce more accurate tentative responses, which can be achieved through reinforcement learning methods such as GRPO. To evaluate their contributions, we apply CRO and GRPO separately to the original MLLMs (e.g., Qwen2.5-VL) and compare their average task performance under the perceptual decoupling framework (details on GRPO are in Appendix C.1). To investigate whether the two objectives are complementary, we also assess a combined setting—denoted as “CRO+GRPO”—where tentative responses and captions are generated by models fine-tuned separately with GRPO and CRO. For comparison, results from models without additional training are labeled as “No Training”. Additionally, we further include VL-Rethinker-7B [47], which is initialized from Qwen2.5-VL-7B and thoroughly fine-tuned using GRPO. We further fine-tune this model with CRO and report the resulting performance as “GRPO+CRO”. All results are visualized in Figure 6.
As shown, both CRO and GRPO improve the performance under the perceptual decoupling framework. More importantly, for all MLLMs, combining both objectives brings notable improvement over the individuals. This suggests that the proposed CRO is complementary to GRPO. Note that we did not report results with GRPO training in Table 1 to highlight the impact of perceptual decoupling.
Figure 7: Pairwise comparison on caption quality.
Evaluating the Quality of Captions.
To see whether the quality of the generated captions are improved, we conduct pairwise comparisons [9, 34] among captions generated by Qwen2.5-VL-3B trained with and without CRO. We randomly sample 1000 captions for each evaluation dataset and use GPT-4o  [^1] https://genai.ust.hk/ as the judge to determine the preferred caption. The judge receives the image, the question alongside two captions, and decides the winner or a tie game. We instruct the judge to choose captions that include more comprehensive and accurate details required to answer the question and exclude any solving process in the captions (the detailed prompt is in Appendix C.2). Furthermore, we alternate the order of the two captions to avoid position bias following [57]. Results are reported in Figure 7. It can be observed that across all datasets, models trained with CRO offer a clear advantage over the original ones. This confirms the quality issues of captions discovered by this paper and highlights the effectiveness of CRO. To ensure robustness of our automated evaluation, we further conduct human assessment through the same pairwise comparison framework with details in Appendix C.2. The results reinforces the quality improvement of the captions.
5 Conclusion
In this paper, we propose the RACRO framework to construct multi-modal reasoning models efficiently. By decoupling multi-modal reasoning with the visual perception and text-only reasoning, RACRO embraces the extraordinary reasoning capabilities of frontier LLMs, while getting rid of the burdensome visual re-alignment. Enhanced with Caption Reward Optimization, precise captions are reinforced to provide abundant visual contexts for better downstream reasoning, effectively scaling towards more advanced LLM reasoners during inference time. We obtain state-of-the-art performance on multiple multi-modal reasoning benchmarks, while remaining computationally efficient.
Appendix
Appendix A Model Configuration for Figure 1
For each RACRO-enhanced model group (e.g., Qwen2.5-VL-3B (RACRO)) in Figure 1, we train the original model (e.g., Qwen2.5-VL-3B, they are shown as hollow squares with different colors in the figure) using both the proposed CRO and the existing GRPO objectives, where the former encourages the MLLM to generate query-conditioned captions with higher quality while the later optimizes it to give better reasoning traces. The details for CRO and GRPO can be found in Sec. 4.1 and Appendix C.1. We then pair the trained MLLM with different LLM reasoners under our RACRO framework. Table 4 shows the configuration for each RACRO-enhanced model group. Specifically, for each model in the group, we report the total model size (B), paired LLMs, data used to conduct GRPO and CRO and the average performance across 7 tasks. Note that the results for Qwen2.5-VL-32B group does not involve GRPO training because we obtain worse results than the original model after GRPO. Results for other MLLMs without RACRO are directly taken from Table 1.
Table 4: Model Configurations for each RACRO-enhanced model group in Figure 1. “-” denotes the corresponding item (e.g., LLM reasoner, GRPO/CRO training) is not applied.
Size (B) LLM GRPO CRO Avg. Performance
Qwen2.5-VL-3B
3 - - - 34.6
7 Qwen3-4B ViRL39K ViRL39K 45.4
11 Qwen3-8B ViRL39K ViRL39K 46.3
35 R1-32B ViRL39K ViRL39K 47.8
Qwen2.5-VL-7B
7 - - - 44.5
11 Qwen3-4B ViRL39K ViRL39K 48.9
15 Qwen3-8B ViRL39K ViRL39K 50.2
39 R1-32B ViRL39K ViRL39K 50.8
@@ Appendix A Model Configuration for Figure 1
32 - - - 52.2
39 R1-7B - ViRL39K 54.2
40 Qwen3-8B - ViRL39K 55.6
64 R1-32B - ViRL39K 56.0
Appendix B Prompt Templates Used in RACRO
In Table 5, we provide an index of the prompt templates used in RACRO.
Table 5: Index of Prompt templates for RACRO.
@@ Appendix B Prompt Templates Used in RACRO
MLLM Tentative response $P_{\text{cot}}$ Figure 10
MLLM Query-conditioned captions $P_{\text{cap}}$ Figure 11
@@
Figure 11: Prompt templates used by the MLLM to obtain the query-conditioned captions.
Appendix C Experimental Details
C.1 Details for GRPO Training
GRPO
@@ C.1 Details for GRPO Training
(Group Relative Policy Optimization) [42] is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs. In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM. For a given input pair $(I,\mathbf{q})$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ rollouts, i.e., $\mathbf{o}\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{cot}}(\mathbf{q}))$. Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$. The objective incorporates a surrogate loss clipped within $[1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
$L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{o}\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{cot}}(\mathbf{q}))}$
@@
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))},\,1-\epsilon,\,1+\epsilon\Bigr{)}\hat{A}_{i}\Biggr{)}\Biggr{]}\textrm{.}$ (3)
@@
The reward $R_{i}$ for the $i$-th rollout is expressed as: $R_{i}=r(y_{\text{gt}},\textbf{o}_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(\textbf{o}_{i})),$ where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise.
Implementation Details.
Similar to CRO, for training data, we adopt ViRL39K [47] and implement GRPO using the verl [43] framework with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$. For each query, eight rollouts are sampled. The clipping constant $\epsilon$ and KL-penalty coefficient $\beta$ are set to 0.2 and $1\mathrm{e}^{-3}$, respectively.
C.2 Details for Pairwise Comparisons

### v2

@@ Results.
Decouple GRPO VPO Cap. penalty Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
70.3 25.8 41.0 57.3 34.4 19.4 46.1 42.0
70.0 40.4 45.2 62.0 39.1 26.3 49.7 47.5
72.7 43.2 50.0 63.3 41.1 28.7 54.1 50.5
76.0 41.5 50.6 62.9 43.1 33.1 57.7 52.2
76.1 43.7 52.2 64.7 45.4 32.7 57.7 53.2
71.2 43.8 48.1 64.6 39.9 32.3 57.9 51.1
74.2 42.5 50.8 62.0 39.4 31.9 56.6 51.1
74.2 29.7 44.8 55.9 41.0 27.7 48.1 45.9
75.0 29.8 42.0 55.8 40.8 23.0 46.3 44.7
4.2 Ablation Studies
In this section, we first investigate the effectiveness of the proposed components, i.e., reasoning-perception decoupling (denoted “Decouple") and VPO. For VPO, we ablate the choices on reward computation during training. Next, we assess the generalization and scalability of RAPID to different LLMs. We use the same training configurations as in Sec. 4.1. Qwen2.5-VL-3B/7B are adopted for ablations due to resource constraints. Unless otherwise specified, we use R1-7B as the default LLM for training (reward computation) and Qwen3-8B for inference.
Reasoning-Perception Decoupling & VPO.
Table 2 presents a detailed ablation study of RAPID’s with the 7B MLLM (see Appendix E.2 for the 3B MLLM), which we analyze by incrementally adding each one to the baseline. Starting from the baseline MLLM (), we first apply perception-reasoning decoupling. This step alone () yields a significant 5.5% average improvement, demonstrating the immediate benefit of leveraging a stronger external LLM (Qwen3-8B) for reasoning. Building on this, we apply GRPO to enhance the MLLM’s perception by optimizing its tentative solutions, which adds another 3.0% to the average score (). We then apply VPO without the caption penalty (), and achieves a further 1.7% gain. Finally, incorporating the caption penalty leads to our full model (), adding another 1.0%. This brings the total improvement from our full VPO method to 2.7% over the model with only GRPO (). This caption penalty is crucial for VPO’s effectiveness, as it prevents reward-hacking where the model might generate solutions instead of the intended captions. Figure 9 confirms this: without the penalty, the ratio of rollouts containing valid captions diminishes rapidly, whereas with the penalty, it remains stable above 95% before 150 steps.
Figure 6: Training dynamics of GRPO and VPO. Performance is evaluated under the perception-reasoning decouple pipeline.
The results also highlight that GRPO and VPO are complementary, whose synergy is evident in two ways: 1) Removing either method from the full decoupled setup ( vs. , and vs. ) results in suboptimal performance, confirming both are necessary. 2) Figure 6 further visualizes this dynamic: after initial gains from GRPO begin to plateau, VPO provides a distinct secondary performance boost. Despite these gains, the decoupling strategy remains the most critical element. An MLLM improved by VPO and GRPO alone () still lags far behind the decoupled version (), underscoring its importance. We also note that VPO does not improve MLLM’s reasoning capabilities on its own ( vs. ); the slight performance drop suggests minor forgetting during sequential training.
Figure 7: Reward hacking without the caption penalty.
Figure 8: Reward dynamics of adopting qcap and cap.
Figure 9: Caption length trend with various LLMs.
While holistic captions (cap+sol) initially outperform query-relevant ones (qcap+sol) as discussed in Section 3.1, this trend reverses after applying VPO (see and ). We hypothesize that this is because qcap is easier to optimize during VPO, as the query guides the MLLM to focus on relevant visual details. Instead, without such guidance, the MLLM struggles to identify pertinent information for cap. This is confirmed by Figure 9, which shows that training rewards for qcap increase steadily, while rewards for cap oscillate without consistent growth.
Choices on Reward Computation.
We study VPO reward computation by varying the reasoning LLM (during training) and its input content. We take the best-performing checkpoint (optimized with GRPO) as the baseline and evaluate under the perception-reasoning decoupled paradigm. First, we keep the input content as qcap+sol and test three LLMs with increasing reasoning capacity: Qwen2.5-7B-Instruct (weak), R1-7B (intermediate), and Qwen3-8B (strong), respectively.
As shown in Table 3, the R1-7B LLM performs best. We hypothesize this is due to a trade-off in reasoning capacity, reflected in caption lengths. Figure 9 shows the caption lengths during training for various LLMs. We speculate that the stronger Qwen3-8B can succeed with more succinct captions,
Table 3: Ablation studies on (a) LLM types and (b) input to LLM for reward computation.
Configuration Q-VL-3B Q-VL-7B
w/o VPO (baseline) 46.1 50.5
(a) LLM Types
Qwen2.5-7B-Instruct 47.5 51.9
R1-Distilled-7B 47.9 53.2
Qwen3-8B 47.1 51.3
(b) LLM Input Contents
qcap+sol 47.4 49.5
qcap 47.9 53.2
thus inadvertently rewarding brevity, while the weaker Qwen2.5-Instruct-7B incentivizes informative captions, causing length to surge. R1-7B strikes an effective balance, making it our default choice.
Next, we examine the choice of input content: using caption alone (qcap) versus using the caption plus a tentative solution (qcap+sol). As in Table 3, using only the caption is superior. Including tentative solutions allows the LLM to take a shortcut during training—relying on the solutions while ignoring captions—which generates a noisy reward signal ineffective for optimizing caption quality. [^5] Notably, this optimal input format differs from that used in the perception-reasoning decoupling stage.
Figure 10: Performance with different LLMs. (Only R1-7B is used in training)
4.3 Generalization and Scaling with Different LLMs.
A practical requirement is that our MLLM, once optimized, should generalize to new, unseen LLMs at inference time without retraining. To test this, we perform VPO on the GRPO-trained MLLM using only R1-7B as the LLM for this optimization step. We then evaluate its performance against the baseline (the MLLM without VPO) by pairing both MLLMs with a diverse range of different LLMs at inference time, as shown in Figure 10. . We have three observations. First, the performance gain from VPO generalizes effectively. The gap between the VPO-trained model (solid curves) and the baseline (dashed curves) is maintained or widened when using stronger LLMs, revealing that the benefit is not confined to the specific LLM used for training. Second, the RAPID’s scalability is evident as absolute performance trends upward when using more capable LLMs, although this improvement is not strictly monotonic with model size. Additionally, among the LLMs tested, Qwen3-8B strikes the best balance between performance and model size, establishing it to be our default choice for the inference stage. Note that the optimal LLMs for training and inference might differ (c.f., Table 3 and Figure 10).
Figure 11: Pairwise comparison on the caption quality with and without VPO.
4.4 Evaluation on Caption Quality
We validate the improved quality of the captions generated by VPO-optimized MLLMs via a pairwise comparisons (Chen et al., 2023; Liu et al., 2024b) using GPT-4o (OpenAI, 2024) as a judge. With 1000 random samples per dataset, GPT-4o compares captions from Qwen2.5-VL-3B trained with and without VPO. The judge is instructed to prefer captions with more comprehensive and accurate details required to answer the question, while excluding any solving process (the prompt is in Appendix E.3). We alternate the caption order to mitigate position bias (Zheng et al., 2023). As shown in Figure 11, the VPO-optimized MLLM’s captions demonstrate a clear advantage across all datasets, highlighting the VPO’s effectiveness (We conduct a case study on caption quality in Appendix G). Moreover, we validate these findings with human assessment in Appendix E.3.
5 Conclusion
This paper proposes RAPID, an efficient method for constructing multi-modal reasoning models. By decoupling visual perception (MLLM) from text-only reasoning (LLM), RAPID leverages the advanced reasoning of frontier LLMs while avoiding burdensome visual re-alignment. Enhanced with Visual Perception Optimization, this method reinforces precise captions to provide rich visual context, improving reasoning and effectively scaling to more advanced LLMs at inference time. Our approach achieves significant accuracy gains on multiple multi-modal reasoning benchmarks while remaining computationally efficient.
Appendix A Model Configuration for Figure 1
For each RAPID-enhanced model group (e.g., Qwen2.5-VL-3B (RAPID)) in Figure 1, we train the original model (e.g., Qwen2.5-VL-3B) using both the proposed VPO and GRPO objectives, where the former encourages the MLLM to generate query-relevant captions with higher quality while the later optimizes it to give better reasoning traces. The details for VPO and GRPO can be found in Sec. 4.1 and Appendix C. We then pair the trained MLLM with different LLMs under our RAPID. Table 4 shows the configuration for each RAPID-enhanced model group. Specifically, for each model in the group, we report the total model size (B), paired LLMs, data used to conduct GRPO and VPO and the average performance across 7 tasks. Note that the results for Qwen2.5-VL-32B group does not involve GRPO training because it has already undergone an RL stage before release (Bai et al., 2025). Results for other MLLMs are consistent with Table 1.
Table 4: Model Configurations for each RAPID-enhanced model group in Figure 1. “-” denotes the corresponding item (e.g., LLM reasoner, GRPO/VPO training) is not applied.
Size (B) LLM GRPO VPO Avg. Performance
@@ Appendix A Model Configuration for Figure 1
7 Qwen3-4B ViRL39K ViRL39K 47.3
11 Qwen3-8B ViRL39K ViRL39K 49.2
17 Qwen3-14B ViRL39K ViRL39K 50.1
@@
11 Qwen3-4B ViRL39K ViRL39K 51.0
15 Qwen3-8B ViRL39K ViRL39K 53.2
21 Qwen3-14B ViRL39K ViRL39K 53.6
@@
36 Qwen3-B - ViRL39K 54.6
40 Qwen3-8B - ViRL39K 55.9
152 GPT-OSS-120B - ViRL39K 57.4
Appendix B Prompt Templates
In Table 5, we provide an index of the prompt templates used in RAPID.
Table 5: Index of Prompt templates for RAPID.
@@ Appendix B Prompt Templates
MLLM Holistic captions $P_{\text{cap}}$ Figure 13
MLLM Query-relevant captions $P_{\text{qcap}}$ Figure 14
MLLM Tentative response $P_{\text{sol}}$ Figure 15
MLLM Caption Penalty - Figure 17
@@
Figure 13: Prompt templates used by the MLLM to obtain the holistic captions.
Figure 14: Prompt templates used by the MLLM to obtain the query-relevant captions.
@@
Figure 17: Prompt templates used by the MLLM for caption panelty.
Appendix C Formulations of GRPO
@@
GRPO (Shao et al., 2024) is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs. In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM. For a given input pair $(I,q)$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ rollouts, i.e., $o\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{sol}}(q))$. Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$. The objective incorporates a surrogate loss clipped within $1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
$L(\theta)={}$ $\mathbb{E}_{(I,q)\sim p_{\mathcal{D}},o\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{sol}}(q))}$
@@
$\Biggl[\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{sol}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{sol}}(q))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{sol}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{sol}}(q))},\,1-\epsilon,\,1+\epsilon\Bigr)\hat{A}_{i}\Biggr)\Biggr]\textrm{.}$ (4)
@@
The reward $R_{i}$ for the $i$-th rollout is expressed as: $R_{i}=r(y_{\text{gt}},o_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(o_{i})),$ where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise.
Appendix D More Analysis
D.1 Analysis on the Training Compute Efficiency of RAPID
RAPID’s decoupled design enables the flexible adoption of recent LLM reasoners, such as Qwen3 (Yang et al., 2025a), without retraining. This raises a critical question: how does our efficient approach compare against models that require full and costly retraining of their visual-language alignment to integrate the latest LLMs?
To investigate this trade-off between performance and computational cost, we compare RAPID with two leading MLLMs also built on the Qwen3-8B LLM: Keye-VL (Team et al., 2025b) and InternVL3.5 (Wang et al., 2025c). Table 6 presents a comparative analysis, reporting average accuracy across seven multi-modal reasoning tasks alongside the training tokens and estimated training FLOPs (calculated as model size × training tokens). The number of training tokens for Keye-VL-8B and InternVL3.5-8B are sourced from their respective technical reports.
As can be seen, although still inferior to the end-to-end methods, RAPID with Qwen2.5-VL-7B can achieve $90.8\%$ of Keye-VL-8B performance with $1250\times$ less training FLOPs, and $88.2\%$ of InternVL3.5-8B performance with $864.2\times$ training cost reduction. Thanks to the remarkable training efficiency, we can adopt larger MLLMs such as Qwen2.5-VL-32B, we achieve comparable ($92.7\%$) performance with InternVL3.5-8B but with $1025\times$ less training FLOPs
Table 6: Training cost comparison. As we did not apply GRPO to Qwen2.5-VL-32B, it consumes less training tokens (100M) than Qwen2.5-VL-7B (550M).
Method AVG Accuracies # Tokens FLOPs (Ratio)
Keye-VL-8B (Team et al., 2025b) 58.6 600B $1500\times$
InternVL3.5-8B (Wang et al., 2025c) 60.3 410B $1025\times$
Qwen2.5-VL-7B w/ RAPID (Qwen3-8B) 53.2 550M $1.2\times$
Qwen2.5-VL-32B w/ RAPID (Qwen3-8B) 55.9 100M $1\times$
D.2 Analysis on the Inference Compute Efficiency of RAPID
To evaluate the inference compute efficiency of RAPID, we estimate the computational cost across the seven evaluation datasets from Table 1. For each dataset, we perform inference on a sample of 100 examples and approximate the compute as model size $\times$ number of generated tokens. For our staged RAPID approach, the total compute is the sum from the perception and reasoning stages, with the compute for each stage calculated using its respective model size. We benchmark these results against top-performing MLLMs from Table 1. Figure 19 plots the resulting average accuracy against inference compute for both RAPID-enhanced models and their baselines.
The analysis reveals two key findings:
  • RAPID achieves a favorable trade-off between accuracy and inference compute. This efficiency is highlighted by specific configurations that demonstrate Pareto-optimality. For instance, Qwen2.5-VL-7B w/ RAPID (R1-7B) provides a better accuracy-to-compute ratio than ReVisual-R1-7B. In another example, Qwen2.5-VL-7B w/ RAPID (Qwen3-8B) outperforms the much larger Qwen2.5-VL-72B while being more computationally efficient.
  • RAPID demonstrates strong scalability with inference compute. The architecture is designed such that allocating more computational resources at inference—for example, by swapping in a more powerful LLM reasoner—consistently yields higher accuracy.
Figure 19: Inference compute versus average accuracy.
Appendix E More Experimental Details
E.1 Evaluation on General Benchmarks
Unlike the specialized math and science reasoning benchmarks, our general evaluation suite is chosen to assess foundational vision-language capabilities. We select MME (Fu et al., 2024a), MMBench-v1.1 (Liu et al., 2024a),
Figure 20: General benchmark Results. (Qwen-2.5-VL-3B)
MM-Vet (Yu et al., 2023), SEED-Image [^6] We evaluated on the “Image” split. (Li et al., 2023), MMstar (Chen et al., 2024), RealWorld-QA (RW-QA) (xAI, 2024), MMT-Bench (MMT) (Ying et al., 2024), and BLINK (Fu et al., 2024b), which cover tasks such as object recognition, text recognition (OCR), spatial awareness and so on. Figure 20 presents the results for Qwen2.5-VL-3B. Similar to the observations for the 7B model (Figure 5), the VPO/GRPO-optimized model performs comparably to the original MLLM (Note we do not report results with Qwen2.5-VL-32B/72B as they show the same observations). This confirms that RAPID preserves general-purpose abilities across different model scales.
E.2 Ablation Study on Qwen2.5-VL-3B
In this section, we extend our ablation study to the smaller Qwen2.5-VL-3B model, with results presented in Table 7. While the results are largely consistent with those from its 7B counterpart (Table 2), a critical difference emerges. The 3B model necessitates an additional GRPO stage () following VPO to restore its reasoning capabilities [^7] We hypothesize that VPO degrades the quality of the intermediate reasoning steps passed to the LLM, an effect not always visible in the final accuracy., which in turn yields considerable accuracy gains. We attribute this requirement to the limited capacity of the 3B model, where optimizing for the VPO task appears to degrade its inherent reasoning performance—a trade-off that is less pronounced in the larger 7B model.
Table 7: Ablation study of different components of RAPID (with Qwen2.5-VL-3B). VPO: VPO without the caption penalty; : using cap+sol for reasoning-perception decoupling. : After VPO, we additionally conduct GRPO to recover its reasoning ability.
Decouple GRPO VPO Cap. penalty Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
65.5 39.1 31.9 59.0 31.1 24.8 48.3 42.8
68.5 40.0 39.2 61.0 35.1 26.9 51.7 46.1
68.5 39.4 43.4 59.9 37.4 27.9 55.3 47.4
69.0 39.7 44.3 60.9 38.6 27.3 55.3 47.9
69.6 40.8 48.6 60.9 39.1 29.3 56.4 49.2
67.0 40.9 44.3 58.0 33.2 28.9 54.4 46.7
68.8 41.0 43.8 59.8 34.8 28.7 54.4 47.3
69.1 26.9 38.2 56.9 34.0 20.0 42.5 41.1
68.8 26.9 39.8 49.4 33.0 21.6 46.3 40.8
E.3 Details for Pairwise Comparisons